OpenStudy (zohaibtarar):

A piece of wire 25mm long and with a rectangular cross section, having a high of 5µm and width of 1mm, has a resistance of 122mΩ. What is the material it is made of? Material ρ (Ωm) at 20 °C Silver 1.59×10−8 Copper 1.68×10−8 Gold 2.44×10−8 Aluminium 2.82×10−8 Tungsten 5.60×10−8 Zinc 5.90×10−8 Nickel 6.99×10−8 Iron 1.0×10−7 Platinum 1.06×10−7 Tin 1.09×10−7 Carbon steel 1.43×10−7 Titanium 4.20×10−7 Stainless steel 6.9×10−7 Nichrome 1.10×10−6

OpenStudy (zohaibtarar):

R =pL/A formula

OpenStudy (shamim):

Can u calculate rho=? From ur given data

OpenStudy (shamim):

Rho=specific resistance

OpenStudy (zohaibtarar):

i m bit confused about the unit conversation etc can you please calculate it? Thank you

OpenStudy (mrnood):

the formula is given clearly above YOU can calculate it just convert all the dimensions to METRES first

OpenStudy (mrnood):

I will convert the FORMULA for you: R =pL/A -> p = RA/L

OpenStudy (mrnood):

122mΩ = 122 *10^-3Ω

OpenStudy (shamim):

1 micrometre=10^-6 m

OpenStudy (shamim):

U know 1 mm=10^-3 m

OpenStudy (shamim):

Can u calculate cross sectional area of ur given wire A=?

OpenStudy (shamim):

A=weidth*height=?

OpenStudy (shamim):

A=1*10^-3m*5*10^-6m=?

OpenStudy (shamim):

I think u hv value of resistance R, cross sectional area A, length of wire L Find out rho=?